LINAC
Shutdown Weeks 1 & 2
- Maintained beam through the low energy Linac to support Neutron Therapy Facility (NTF) first two weeks of shutdown
- I- Source Column inspection found flaking/cracked ceramics. Cleaning and polishing in progress
- LRF5 special beam run to dump to test new cavity loop configuration was promising
- Converted stations to new heliax for low level RF upgrade. Phased in new system
- Conducted low level RF studies and code tests
- Prepared for planed 6/22 power outage and made good recovery of low energy RF stations to resume NTF support in second week of shutdown
- Began maintenance on low energy (LE) stations 4 & 5 second week of shutdown
- Began high energy (HE) station maintenance second week of shutdown.
- Began LE Station 4&5 480V breaker panel and wiring upgrade in week 2
- Began HE klystron water system maintenance in second week.
- Worked on PFN SCR water lines
- General cleaning and maintenance of LE and HE stations well underway
- MI8 Booster LCW reconfiguration to get off of MI LCW system almost completed
- Cleared 18 tunnel penetrations for cable pulls
- RF cavity 14 pulled to replace bottom tuner. Work almost completed
- East Gallery valved out to allow return pipe upgrade work
Booster
Shutdown Weeks 1 & 2
- Correction element upgrade
- 19 out of 36 magnets installed
- Need to complete magnet vacuum interface connections and pump down
- Discrepancy found in alignment of magnetic center (Stretch wire) and physical center of magnets. Will align to physical center. Analyses shows this will be fine
- GMPS upgrade in progress. Pads stripped of old Brentford cubicles and oil transformers
- MP02/MP02 Polarity change in progress
- HLRF cleaning and maintenance in progress
- New L12 kicker beam tube and girder in place

Antiproton Source
Accumulator 4-8GHz DP Kicker Upgrade: ~30% complete
- All beam line components are in place. Final alignment next week(?), then pump down & leak check. Bake will follow. A40 bake is already underway and will finish on Saturday
Debuncher Septum Magnet Replacement: 0% complete
- Will start unstacking shielding steel this weekend. New magnet is in the tunnel
IB7 (AP2 dipole magnet) Replacement: 0% complete
- This work will probably begin in August. The LCW has been turned off to allow magnet and floor time to dry out before work begins
Maintenance work on various RF systems.
Installing beam pipe & vacuum value in preparation for resurrection of anti-hydrogen experiment. This might add a bake to our plans.
Various LCW leaks being addressed.
A60 cryo warm up (LN2) for cryostat vacuum system maintenance work and possible bake.
Main Injector
MI10 area: On Schedule
- Lead blankets have been installed
- Magnets have been wrapped
- RR magnets stand installed
- All but two catch pans installed
- Excavation is close to the 5 foot point
MI40 area: On Schedule
- Lead blankets have been installed
- Magnets have been wrapped
- Catch pans are installed
- Excavation to the tunnel is complete
- Boring to begin next week
Instrumentation: On Schedule
- Replaced micro channel plates for all three IPMs
- Replaced Vertical Flying Wire
RF: On Schedule
PS: On Schedule
LCW: Ahead of Schedule
- Finished with all non-hot spots, working on A1/P1 lines
MI 308 Vacuum leak: Not Scheduled but On Schedule
MI-8 Multiwires:
- One MW being vacuum certified
- Have parts to assemble six more
Recycler
Recycler
- Pot RR129 (i.e. MI-10 injection kicker)
- Move 6 magnets
- HLS installation at MI-10 (possibly)
Ecool
Biggest job: replace charging circuitry ( chain, wheels, brushes…)
- Flush Water/Glycol skid (and magnets)
Decided (late) that it was safer to vent and purge with N 2 the two sections of the Recycler that are in the vicinity of the MI tunnel digging/drilling/coring work ~240 m of the ring (7.5%) under N 2 for 1-2 months
Tevatron
Roll Measurements
- Rolls on 954 magnets were measured in the first week of the shutdown
- 11 students from CDF and D0 and 4 Operators did the measurements
- 88 magnets with rolls greater than ± 1 milli radian were found
- Of these 6 magnets had rolls greater than 2 milli radian
- There is data for every shutdown from 03. The data is at http://home.fnal.gov/~volk/TeVrollmeas/
Other Measurements
- As found low beta quads and separators at D0 and B0 done
- Level run through tie rods around TeV done
- Cold leaks investigated in A and F sector
- A3 house warmed to room temp to fix cold leaks, bad spool at A34 removed
- Started un rolling TeV dipoles and quads
- Removed B-2 dog leg to make room for vertical Crystal Collimator at E0
- FESS re lamping entire ring mostly done
- Working through TeV Fix list
NuMI
Shutdown Weeks 1 & 2
- Removed Shielding plug & installed Crane electronics
- Change water in Horn 1 & 2 plus Target RAW skids
- Replace water flow meters in Target Hall
- Installed Mass Gas Flow Meters on RAW Surge Tanks for Horn 1 & 2
- Relocated thermocouple electronics from RAW to Power Supply Room
- Complete initial Rad Surveys in the Target Pile
- Replaced leaking LCW valve on Horn 1
- Installed all Target Hall & Work Cell cameras & upgraded Work Cell hydraulics wiring
- Replaced & shielded Target Pile Humidity Monitors
- Change out Horn 1 DI Bottles
- Completed initial investigation of Horn 2 water leak
- As-found survey completed of the Pre-Target and Target Hall
- Completed as-found survey of the Work Cell rails
- Completed Underground Facility Inspection by Hanson Engineering including crane rail
- Floor drain water samples taken from Pre-Target to Absorber Hall
- All underground floor drains cleaned
- Removed Rowald chiller from mechanical room
- Absorber Hall Dehumidification work started
- Electrical utility relocation completed for Hadron Monitor Replacement
- As- Found survey of EPB Dipole HV101 string completed
- Work started on HV101-1, 101-2, 101-6 shielding removed and PEEK installation started on 101-1
- Profile Monitor work re-scheduled for Week 8
Cryogenic System
- Support AD/MSD for A3 and D3 warm-up/isolation for repairs
- Ring wide vacuum pump maintenance, repairs, and pumpdown of various LN2 components to assure stable Tevatron LN2 Mode Operations.
- Pbar Accumulator LN2 tanks disconnected per Pbar request.
- 6 wet engine flywheels replaced, 2 wet engine motors replaced, 1 wet engine overhauled
- Support FESS work on various TEV pond work re Mycom compressor operations; C0 water pump start/stop control button failed (FESS procuring replacement); E0 compressor Yard breaker failed (FESS replaced breaker); A0 water pump remote start relay failed (replaced); B0 water pump high vibration due to cracked housing (FESS replacing it).
- Ongoing FRIG/Compr controls/instrumentation maintenance/repairs.
- CHL at room temperature for maintenance. LN2 supply to Tevatron relying on Linde deliveries (3-4 per day).
- Purifying Tevatron via CHL Main Helium Purifier; ready for cooldown July 1 (just kidding)
Controls
So Far:
- Networking
- Backbone switch upgrade
- Main Injector Subnet Address change
- Tevatron Zero Bldg. switch upgrades (2/3 complete)
- Switch IOS updates underway
- System Fileserver (CENTRA & CLXSRV) OS Upgrades
- Console node OS Updates
- General Maintenance (filters, fans, power supplies, etc.)
Upcoming Week:
- AD General Network Switch Upgrade (Saturday afternoon – 4hr.)
- Affects AD desktop/laptop systems
- No email, internet, network drive, printer or PC consoles
- UPS Maintenance - Network & TLG (6/30 & 7/1)
- General Maintenance
FESS
Electrical and High Voltage
- MI - Lighting completed except for ballasts & stairwells
- TeV – Lighting re-lamp WIP for next 2 – 3 weeks ON 4 – 12 SHIFT
- TeV – FDR 46A[AO/CO/EO] re-scheduled to 6/23[completed] from 8/6 due to E0 2500a breaker malfunction/replacement for He compressors
- AD/Bob M schedule – all work completed for High Voltage
- FINDINGS – loose connections on pond pump feeders, CDF/B0 HV cable required re-termination
Mechanical
- TeV - B0 [A] 75hp pond pump failed [cracked pump housing/base]
- TeV & MI - strainers cleaned – some ZMs
- TeV - pond intake screens cleaning WIP with divers – no major findings, silt levels ok
- TeV - sump pumps & tunnel supply/exhaust fans completed
- MI – pond intake lines camera inspections completed
- SWYD enclosure B & transfer hall sump pumps completed
- Meson area mechanical PMs completed
- Findings – MI 10 Sump pump repair completed
- Next week – Booster Ring sump pumps, TeV pond pumps, MI tunnel make up air fans
HVAC
- AP 0 & A0 kicker - completed
- Meson area – completed
- No major findings
- Next week – MI tunnel de-humidifiers.
Other Work and Comments
- BGE/BTE sink hole due to sump discharge line break – scheduled for Sat/Sun 6/27-28
- GCC outage on 6/22 completed
- C-4 pump house wet well cleaned by divers
- NUMI MI 65 sump pump replaced
- Casey’s Pond tentative shut down for 7/7 being reviewed regarding ICW water requirements & the need for a base line water supply. At this time the C4 pump house can only accommodate 1500 – 2000 gpm. A 1 hr shut down of Casey’s is planned on 7/6 to determine action plan with possible partial use of Casey’s
Safety
Shutdown Injuries
- One minor finger laceration – First Aid Case
- Investigating a knee injury
Radiation Safety
- All individuals working in the enclosures need to report pocket dosimeter readings
- Presently compiling a list of people we know are working in the enclosures that we are not receiving reports from
- MuCool Test Area (MTA) crane work to begin next week
Interlock Group
- Oxygen cell replacements - 85% complete
- Safety System Interface Unit capacitor replacements - 30% complete
- Single Line Electrical Drawing Verifications - 15% complete
Interlock Work
- Oxygen monitor cell replacements - 85% complete
Safety System Interface Unit capacitor replacements - 30% complete
Single Line Electrical Drawing Panel Verifications - 15% complete
Continue using the Take Five principles Before, During, and After work activities to:
- Avoid Injuries
- Reduce Environmental Impacts
- http://www.fnal.gov/takefive/
|